### v1

# XL-Instruct: Synthetic Data for Cross-Lingual Open-Ended Generation

Abstract
Cross-lingual open-ended generation—i.e. generating responses in a desired language different from that of the user’s query—is an important yet understudied problem. We introduce XL-AlpacaEval, a new benchmark for evaluating cross-lingual generation capabilities in Large Language Models (LLMs), and propose XL-Instruct, a high-quality synthetic data generation method. Fine-tuning with just 8K XL-Instruct-generated instructions significantly improves model performance, increasing the win rate against GPT-4o-Mini from 7.4% to 21.5%, and improving on several fine-grained quality metrics. Additionally, models fine-tuned on XL-Instruct exhibit strong zero-shot transfer to both English-only and multilingual generation tasks. Given its consistent gains across the board, we strongly recommend incorporating XL-Instruct in the post-training pipeline of future multilingual LLMs. To facilitate further research, we will publicly and freely release the XL-Instruct and XL-AlpacaEval datasets, which constitute two of the few cross-lingual resources currently available in the literature.
1 Introduction
Cross-lingual generation is the task of understanding a query in a given source language and generating a response in a different target language. This task has assumed greater relevance in the recent development of Large Language Models (LLMs) with multilingual capabilities. Marchisio et al. (2024) noted its usefulness for a) companies that serve such LLMs across dozens of languages but optimizing a prompt for each input language is inefficient in practice, and b) when a user needs a generation in a language they do not speak. The conventional two-step ‘Reason then Translate’ approach to cross-lingual generation (Huang et al., 2023; Qin et al., 2023; Li et al., 2024b) can be problematic due to the noisy nature of machine translation (MT), which may lead to information loss or an unnatural-sounding response. It is also wasteful of inference time and cost, since the intermediary English response is thrown away once the desired cross-lingual output is obtained.
Despite its relevance, the adaptation of LLMs for cross-lingual generation is insufficiently studied. In the absence of high-quality instruction datasets and evaluation benchmarks, the cross-lingual abilities of LLMs remain emergent, yet inadequately assessed. In this work, we address the data deficiency for the cross-lingual generation task. We introduce XL-AlpacaEval, a cross-lingual evaluation dataset sourced from AlpacaEval (Li et al., 2023b), and observe poor off-the-shelf performance for most open multilingual LLMs. As a solution, we propose XL-Instruct, a synthetic data generation technique to create high-quality cross-lingual data at scale (illustrated in Figure 1) and show that fine-tuning with XL-Instruct significantly and consistently boosts cross-lingual performance across a range of base and instruction-tuned LLMs. Beyond cross-lingual capabilities, we demonstrate how this technique also shows strong zero-shot transfer performance in monolingual generation scenarios, both in English and in other languages. We will publicly release the XL-AlpacaEval benchmark and the XL-Instruct dataset, hoping to facilitate research in the cross-lingual LLM domain, which currently lacks sufficient resources for both evaluation and post-training.
We seek to answer the following Research Questions in this work:
  • RQ1: How good are SOTA multilingual LLMs on cross-lingual generation off-the-shelf? (§3)
  • RQ2: How well does XL-Instruct improve cross-lingual generation performance? (§3)
  • RQ3: How much does XL-Instruct transfer to zero-shot monolingual generation scenarios, both in English and non-English settings? (§6)
Figure 1: The XL-Instruct pipeline: 1) instruction generation from seed English data; 2) data refinement; 3) response translation into non-English; 4) data filtering, with more details in Section  4.
2 Related Work
Cross-Lingual Explorations in LLMs
Most of the current work on cross-lingual techniques in LLMs has focused primarily on prompting strategies . The popular approach is to leverage the extensive knowledge and superior reasoning capabilities of LLMs in high-resourced languages (like English) to improve performance in relatively lower-resourced ones (Qin et al., 2023; Huang et al., 2023; Singh et al., 2024; Wang et al., 2025). PLUG (Zhang et al., 2024) also shares this motivation, but fine-tunes an LLM to respond to a non-English question by first answering cross-lingually in English, followed by a translation to the target language. Extensions to this cross-lingual prompting paradigm have also emerged, such as X-InSTA (Tanwar et al., 2023) which uses a semantic encoder to select relevant cross-lingual examples, while SITR (Li et al., 2024b) employs self-reflection and iterative refinement to improve cross-lingual summarization. However, no prior study has approached cross-lingual open-ended generation as the primary training objective.
Synthetic Data Explorations
Previous studies on the creation of synthetic data for post-training LLMs has often been limited to English. Self-Instruct (Wang et al., 2023a) and Unnatural Instructions (Honovich et al., 2023) were among the earliest works to demonstrate how synthetic data pipelines could be used to instruction-tune an LLM. Some other efforts looked at the generation of diversified and high-quality data. Tülu 3 (Lambert et al., 2024), for instance, used persona-driven prompting to yield diverse synthetic data (Ge et al., 2024), while Llama 3 (Dubey et al., 2024) leveraged skill-specific experts as teacher models to generate data for coding, math, multilinguality, etc. To enable multilingual support, MT is often used to extend English resources to other languages (Muennighoff et al., 2023; Lai et al., 2023; Ranaldi & Pucci, 2023; Chen et al., 2024). Given the English resources are often model outputs themselves (eg. of chatGPT), training on translations of these can limit models’ exposure to diversity.
Reverse Instruction
There is a subset of data synthesis approaches called ‘reverse instruction’ methods which propose generating instructions from seed data, and then using the original seed data as responses to these instructions. Our work follows this trend of approaches. Initial works in this space Li et al. (2023a); Wang et al. (2023b) presented a two-step procedure which can be done iteratively: 1) fine-tuning a model to perform instruction generation, followed by 2) heuristic-based filtering to keep high-quality synthetic data. Additionally, Chen et al. (2023) also suggested “instruction wrapping” to refine response quality before fine-tuning the reverse instruction model. LongForm (Köksal et al., 2023) bypassed the fine-tuning stage in these works and leveraged a “teacher” LLM (InstructGPT) to generate such instructions off-the-shelf, yielding significant improvements in English text generation tasks. MURI (Köksal et al., 2024) and X-Instruction (Li et al., 2024a) extend LongForm to multilingual generation. The former back-translates to English, generates reverse instructions, and then forward-translates to low-resource languages. The latter bypasses back-translation to English and queries the teacher LLM in the low-resource language directly, potentially exposing the synthetic data to quality issues. The focus of these works is on improving ‘monolingual’ generation performance, ie. where query and response are in the same (non-English) language. Finally, Iyer et al. (2024a) and Iyer et al. (2024b) use similar strategies to create low-resource cross-lingual data for improving the translation performance of LLMs.
Unlike these previous works, our primary goal is to contribute data resources for cross-lingual open-ended generation, which includes a synthetic dataset where the instruction and response are in different languages, as well as a cross-lingual evaluation benchmark. [^1] To the best of our knowledge, we are the first work to propose a cross-lingual open-ended generation benchmark, and our synthetic training dataset is one of the few publicly available. Our experiments (see Table 9) show that it is of much higher quality than the closest prior work, X-Instruction (Li et al., 2024a). We intend to release the XL-Instruct dataset under permissive open source license.
3 XL-AlpacaEval: A Cross-Lingual Evaluation Benchmark
Dataset
To benchmark cross-lingual open-ended generation, we create XL-AlpacaEval, which is adapted from the original AlpacaEval dataset (Li et al., 2023b). AlpacaEval contains 805 multi-domain prompts sampled from various test sets (Dubois et al., 2024), which includes OpenAssistant (Köpf et al., 2024), Koala (Geng et al., 2023), Vicuna (Chiang et al., 2023), Self-Instruct (Wang et al., 2023a) and Anthropic’s Helpfulness test set (Bai et al., 2022). Evaluation is carried out through the LLM-as-a-judge (Zheng et al., 2023) approach, wherein performance is measured by comparing the win rates of model outputs against reference (another model) outputs, as judged by an evaluator model.
To create XL-AlpacaEval, we first manually examine the AlpacaEval datasets and filter out prompts that are tailored towards eliciting responses in English. For example, questions about correcting grammar in an English sentence cannot be answered cross-lingually (refer Section A.1 for a detailed justification and list of prompts that were excluded). The filtered test set consists of 797 prompts. Next, we add cross-lingual generation instructions (such as “Answer in lang language”) to each of these prompts. We randomly sample these instructions from a list of templates (refer Section A.2 for a complete list) and create an evaluation set for 8 languages: 2 High-Resource EU including German (deu) and Portuguese (por), 2 Medium-Resource EU including Hungarian (hun) and Lithuanian (lit), 2 Low-Resource EU including Irish (gle) and Maltese (mlt), and 2 non-EU languages including Chinese (zho), and Hindi (hin). We also focus on evaluating the En-X direction in this work, as it might be the more common use case due to the prevalence of English text on the internet. It would be straightforward to extend this to other languages and pairs—one would simply need to run a script to append the cross-lingual templated instructions to our filtered test set.
Evaluation
While the original implementation used GPT-4-Turbo as both reference and evaluator models, we use GPT-4o Mini for reference and GPT-4o as the judge, given the SOTA multilingual capabilities of the GPT-4o models. Our choice of using GPT-4o Mini as the reference model is motivated by two reasons: i) we experiment with ~7B-9B LLMs in this work, making the Mini model a suitable baseline; and ii) using different reference and evaluator models, with the more capable one assigned the role of the latter, should mitigate self-preference bias of models (Wataoka et al., 2024). Finally, GPT-4o has also been shown to obtain SOTA pairwise correlations with human preferences in multilingual chat scenarios (Gureja et al., 2024; Son et al., 2024) — making it suitable for our task.
Models
To evaluate the off-the-shelf cross-lingual capabilities of current multilingual LLMs, we benchmark several strong open-weight models in the 7B-9B parameters range: Aya Expanse 8B (Dang et al., 2024), Llama 3.1 8B Instruct (Dubey et al., 2024), Gemma 2 9B Instruct (Team et al., 2024), Qwen 2.5 7B Instruct (Yang et al., 2024), EuroLLM 9B Instruct (Martins et al., 2024), Aya 23 8B (Aryabumi et al., 2024) and Salamandra 7B Instruct (Gonzalez-Agirre et al., 2025). Inference of all models in this work is performed using the AlpacaEval repository (Li et al., 2023b), with the default decoding settings, i.e. temperature is 0.7, maximum tokens are set to 2048, and all models are loaded in bfloat16.
Results (Zero-Shot)
We show our benchmark scores in Table 1. Aya Expanse leads the table, achieving a  60% win rate against GPT-4o Mini for the four languages from X-AlpacaEval which it supports (por, deu, zho, hin). While it was trained on significant synthetic data itself using multilingual experts (Dang et al., 2024), it is not clear from the reportage whether its superiority is due to explicit cross-lingual tuning or implicit transfer. For the remaining languages, Llama 3.1 and Gemma 2 perform comparably with win rates ranging between 10% and 30%. We make two critical observations here. Firstly, except for Aya Expanse, most open LLMs trail significantly behind GPT-4o-Mini in cross-lingual generation, leaving significant room for improvement. Secondly, the performance varies widely based on the resourcefulness of the language. While Aya Expanse, Llama 3.1 and Gemma achieve win rates of 40% or higher for high-resource languages like por, deu and zho, performance drops to 20-30% for medium-resourced languages (hun, lit, hin) and even worse for lower-resourced languages like gle and mlt (10% or less). This suggests that we need a reliable pipeline to scalably create synthetic data across varied languages, without substantial quality drops for the lower-resourced ones, to ensure more consistent performances (see Table 6).
Model Avg High-Res EU Med-Res EU Low-Res EU Non-EU
por deu hun lit gle mlt zho hin
Salamandra 7B Instruct 6.44 8.64 8.27 5.08 9.51 5.63 4.95 5.24 4.23
Aya 23 8B 8.85 17.04 15.04 2.07 2.22 2.45 1.92 9.46 20.57
EuroLLM 9B Instruct 12.70 18.94 16.49 8.66 16.57 9.37 8.51 14.82 8.23
Qwen 2.5 7B Instruct 16.73 30.88 16.35 6.82 14.68 7.17 3.69 44.63 9.59
Gemma 2 9B IT 23.29 35.42 32.08 19.80 27.28 10.09 10.03 28.12 23.50
Llama 3.1 8B Instruct 24.36 40.28 35.72 23.07 20.74 13.20 8.47 31.21 22.22
Aya Expanse 8B 35.67 62.75 60.27 8.62 19.54 10.43 9.51 57.22 56.99
Table 1: Win rates against GPT-4o (Mini) on XL-AlpacaEval, as judged by GPT-4o. Languages are denoted by their ISO 639-3 codes.
Model Avg por deu hun lit gle mlt zho hin
Salamandra 7B Instruct 4.45 3.32 2.47 2.16 3.71 7.49 8.00 6.09 2.37
Aya 23 8B 1.28 1.12 -1.78 -8.62 4.58 4.52 11.86 3.11 -4.59
EuroLLM 9B Instruct 5.26 -1.57 -0.83 5.50 5.14 18.66 6.83 11.85 -3.54
Qwen 2.5 7B Instruct -1.25 -20.01 2.92 1.59 2.91 3.62 4.24 3.80 -9.08
Gemma 2 9B IT -4.73 -11.00 -12.66 -4.10 -2.58 2.67 -0.37 6.54 -16.37
Llama 3.1 8B Instruct -10.55 -23.84 -18.01 -7.96 -8.14 -1.75 -2.69 -0.08 -21.92
Aya Expanse 8B -20.53 -39.60 -39.25 -36.13 -0.51 -1.18 -2.50 -1.78 -43.29
Table 2: Performance changes on using Reason-then-Translate: scores represent differences against win rates from Table 1. Strong positive improvements are shaded.
Baseline: Reason-then-Translate
Previous work had proposed prompting LLMs to reason first in a high-resource language such as English and then translating into the target language (Qin et al., 2023; Huang et al., 2023; Wang et al., 2025). We call this approach ‘reason-then-translate’ and report results in Table 2. The outcomes are mixed: stronger multilingual models like Aya Expanse, Llama, and Gemma suffer significant performance drops. Manual inspection reveals these 7B models occasionally produce empty outputs, likely due to difficulty consistently following complex instructions. We note that previous studies have only explored this strategy with larger ( 70B) models. In contrast, weaker LLMs like EuroLLM and Salamandra, fine-tuned mainly on English reasoning and MT data, can leverage this two-step approach to slightly improve their initially poor performance. Overall, our results indicate that improving cross-lingual generation performance is challenging and may not be easily resolved through prompting strategies alone.
4 The XL-Instruct Pipeline
To address this gap, we introduce the XL-Instruct pipeline (illustrated in Figure 1), designed to create cross-lingual synthetic instructions from a given seed corpus. At this point, we highlight two important considerations. First, unlike related work (Li et al., 2024a), we select an English seed corpus for data generation. Our rationale here is to optimize the quality of synthetic data produced by teacher LLMs, as we hypothesize that these models should be more proficient at generating diverse and high-quality synthetic data in English. MT is employed only in the final stages, thereby minimizing noise propagation. Second, we exclusively utilize open-weight models with permissible licenses to generate synthetic data, aligning with our objective of releasing a fully permissible, open-source dataset.
We outline the four stages of the XL-Instruct pipeline in detail below:
  1. 1.
    Stage 1: Reverse Instructions Given a passage from our seed data, we ask a teacher LLM to generate an instruction for which this passage would be a valid response.
  2. 2.
    Stage 2: Refinement: Then, we ask the teacher to reword the question and response such that it follows four manually-defined criteria.
  3. 3.
    Stage 3: Response Translation Next, we translate the refined response, leveraging one or more translation LLMs, to the target language.
  4. 4.
    Stage 4: Filtering To ensure we use the highest quality targets, we use MT Quality Estimation (QE) models to select the best response translations in our dataset.
We conduct Supervised Fine-Tuning (SFT) on this synthetic dataset. We detail the minutiae of each stage below.
4.1 Stage 1: Question Generation
First, we sample an English passage from our seed corpus, CulturaX (Nguyen et al., 2024). Then, we ask a teacher LLM - here, Qwen 2.5 72B (Yang et al., 2024) - to produce an instruction for which the sampled sentence would be a valid response. Prompting in English allows us to leverage the teacher model directly without requiring the additional fine-tuning employed in previous work (Li et al., 2024a). To convert the LLM response to a cross-lingual instruction, we add a prompt to “Respond in {lang}” where {lang} is the target language of interest. This creates a potential synthetic instruction for SFT — the desired output for which would be the response translated to this language (see Section 4.3).
4.2 Stage 2: Refinement
Next, inspired by Self-Refine (Madaan et al., 2023), we use the teacher LLM (again, Qwen 2.5 72B) to refine the question-response pair further. Based on the most commonly occurring errors observed from manual inspection, we define four goals for the refinement process:
  1. 1.
    Question Self-Sufficiency: The question should be clear, unambiguous, and should not require any additional information/context to produce the given response.
  2. 2.
    Response Naturalness: The response should be ‘natural-sounding’ as an LLM output — in terms of fluency, neutrality and objectivity, and consistency with the tone and style of LLM-generated responses.
  3. 3.
    Response Precision: The response should be topically relevant, factually accurate and should directly answer the question. This can be thought of as analogous to precision since it tries to assess how much of the information contained in the response is actually relevant, necessary and true.
  4. 4.
    Response Informativeness: The response should be informative and helpful, and should contain enough justification and explanation to make it useful to an end user. This is similar to recall, as it evaluates how much of the relevant and useful information for the response is actually provided.
We provide all four criteria and their explanations in a prompt and ask the teacher to refine the (question, response) pair. We also instruct the model to ensure the reworded response is grounded in the original one, and request it to not add any of its own knowledge — since we want to avoid excessive teacher distillation and ensure our synthetic data is grounded in the seed corpus to optimize for diversity.
4.3 Stage 3: Response Translation
Next, the response must be machine translated into the target language. Since document-level MT by open LLMs is currently unreliable due to limited exploration, scarce datasets, and hallucination risks, we instead use sentence-level translation. We sentence-split using Segment Any Text (Frohmann et al., 2024) and generate translations in one of two ways:
  1. 1.
    Naive: In the vanilla case, we simply prompt an LLM for the translation.
  2. 2.
    Best-of-k: We obtain $k$ translations from $k$ different LLMs for each sentence, and choose the one with the best QE score as the final translation.
For QE, we use the WMT’23 CometKiwi-XL model (Rei et al., 2023), which obtained state-of-the-art (SOTA) scores in the WMT 2023 QE Shared Task (Blain et al., 2023). For MT, we use EuroLLM 9B Instruct (Martins et al., 2024) in the ‘naive’ case due to its strong MT capabilities, while in the ‘Best-of-k’ setting, we set $k=3$ and sample among EuroLLM 9B Instruct, Mistral Small 24B Instruct (Mistral AI Team, 2025) and Gemma2 27B Instruct (Team et al., 2024). Finally, we add a postprocessing step that involves creating the response translation by substituting each sentence-level translation in-place in the original response. This helps us retain formatting (like paragraph separators, bullet points etc.) that is critical to response quality.
4.4 Stage 4: Filtering
Finally, to ensure that we select high-quality targets during fine-tuning, we compute sentence-level QE scores using the WMT’23 CometKiwi-XL model, by comparing each source sentence in a given response and its translation. We average these QE scores across the entire passage to obtain the final passage-level score. Then, we sort all responses in descending order and filter the last 20% of dataset – creating a final dataset of about 32K instructions. We provide the prompts used in each stage at https://tinyurl.com/xli-prompts (to be uploaded to GitHub on acceptance)
5 Experiments on XL-AlpacaEval: Boosting Cross-Lingual Generation
(a) Base LLMs
(b) Instruction-Tuned LLMs
Figure 2: Performance on XL-AlpacaEval after SFT with XL-Instruct data of varying sizes. Y-axis scores reflect win rates against GPT-4o-Mini, averaged across 8 languages, with GPT-4o as the judge. X-axis instruction counts are shown on a log scale.
Models
We conduct SFT of two base (EuroLLM 9B, Qwen2.5 7B) and three instruction-tuned (EuroLLM 9B Instruct, Qwen2.5 7B Instruct and Aya Expanse 8B) models. We choose EuroLLM and Qwen since they relatively underperform on the XL-AlpacaEval benchmark (Table 1), leaving significant scope for improvement. We also experiment with Aya Expanse since it leads the benchmark, and we are interested in seeing how much further it could be improved. Unfortunately, Aya Expanse does not have a base model released, so we are unable to experiment with it.
Experimental Setting
We fine-tune all models for 1 epoch using low-rank adaptation (LoRA, Hu et al., 2022) with rank 8 matrices applied to query and value projections. Training used a cosine learning rate scheduler with a peak learning rate of 1e-4 and 3% warmup steps. We employed bf16 mixed-precision training with batch size 8, and fix the random seed as 1 for reproducibility.
Main Results
In Figure 2, we show the improved average win rates obtained by fine-tuning with various amounts of XL-Instruct data. For base LLMs, performance steadily improves with data scale until 8K training examples, at which point we observe saturation. Interestingly, EuroLLM consistently outperforms Qwen. The reason for this is revealed from the language-specific scores (Table 6) – the largest improvements are yielded for the languages the model is pre-trained on. For EuroLLM, which includes all 8 languages in its pre-training, there are significant, consistent, and relatively uniform gains across all languages, including the least-resourced ones. Qwen, which chiefly supports high-resource languages like Chinese, Portuguese and German, gains the most for these pairs but shows relatively smaller improvements for the others. This suggests that while post-training with XL-Instruct can yield stable improvements across multiple languages, multilingual pre-training is crucial for best performance.
For instruction-tuned models, the saturation occurs much sooner—at around 2K instructions for EuroLLM-9B-Instruct and, only 32 instructions for the Qwen and Aya Expanse models! This can be explained by the fact that the latter two models have also undergone post-training with Preference Optimization—therefore, task-specific SFT at scale might lead to overfitting and deteriorated performance. EuroLLM has only been exposed to SFT data, and can therefore be trained for longer. In all cases, though, one observes consistent and significant improvements across all languages—with only 32 examples, we can push the win rates of Aya Expanse 8B to $\sim 65\%$ for its supported languages of Portuguese, German and Chinese (Table 6).
Finally, we show in the Appendix (Table 7) how XL-Instruct can also boost zero-shot cross-lingual performance, ie. even for languages not included during SFT.
Fine-Grained Evaluation
Beyond win rates that depend solely on pairwise comparisons, we are also interested in evaluating how well the produced cross-lingual generations improve on an absolute scale, on human-desired criteria. To achieve this, we take inspiration from recent works that define customised, task-specific metrics and use LLM-as-a-judge for producing scores on a Likert scale – achieving significant correlations with human ratings on evaluation of tasks like summarization (Liu et al., 2023), retrieval (Upadhyay et al., 2024), story generation (Chiang & Lee, 2023), Machine Translation (Kocmi & Federmann, 2023b; a) and open-ended generation (Kim et al., 2023). In particular, Kim et al. (2023) showed that using clearly defined rubrics can result in Spearman correlations up to 0.87 with human preferences for open-ended generation. Inspired by this, we propose four criteria pertinent to the task of cross-lingual generation: Objectivity, Naturalness, Informativeness, and Precision. We define detailed rubrics for each metric and provide well-defined criteria for mapping output quality to scores on a scale of 1-5. We include these rubrics in the context of a prompt, and ask GPT-4o to score cross-lingual generations of EuroLLM 9B, EuroLLM 9B Instruct and EuroLLM 9B XL-Instruct (the best model from Figure 2(a), which is fine-tuned with LoRA on 8K examples). We provide detailed evaluation prompts and rubrics at https://tinyurl.com/xl-gen-eval (to be uploaded to GitHub on acceptance).
Figure 3: Evaluations on fine-grained quality metrics by GPT-4o on a scale of 1-5. For EuroLLM 9B XL-Instruct, we choose the best model from Figure 2(a). Despite being LoRA fine-tuned on only 8K instructions, it performs comparably or slightly better than EuroLLM 9B Instruct which has been fully fine-tuned on a 2M examples dataset. These scores are also tabulated in Table 8.
We show our results in Figure 3, which provides the macro-averaged scores for each criterion per model. As one would expect, the untuned EuroLLM 9B base model achieves the worst scores on all metrics, with the EuroLLM-9B-Instruct model performing substantially better. We note that the XL-Instruct model performs comparably to or marginally better than the EuroLLM-9B-Instruct model. This result is particularly impressive given the XL-Instruct model was trained using LoRA fine-tuning on only 8K synthetic samples, whereas the EuroLLM-9B-Instruct model was fully fine-tuned on 2M human and synthetically-generated examples. These results clearly demonstrate the effectiveness and high quality of the XL-Instruct dataset.
6 Experiments on m-AlpacaEval: Exploring Zero-Shot Transfer
Model Avg zho deu hin hun gle lit mlt por eng
EuroLLM 9B 0.73 1.19 1.47 0.70 0.14 0.14 0.65 0.31 1.25 35.59
+XL-Instruct (best, 8K) 6.10 10.77 11.40 4.47 2.53 3.60 3.77 2.49 9.76 51.35
EuroLLM 9B Instruct 8.94 13.38 11.99 8.13 4.81 5.65 6.68 6.78 14.12 55.58
+XL-Instruct (best, 8K) 15.55 19.57 18.30 13.03 10.38 14.12 16.76 14.13 18.11 59.44
Qwen 2.5 7B 2.04 10.40 1.52 0.98 0.24 0.03 0.45 0.29 2.39 46.93
+XL-Instruct (best, 8K) 5.66 20.43 9.53 2.23 0.65 0.18 1.60 0.29 10.33 55.92
Qwen 2.5 7B Instruct 11.47 45.29 10.53 5.71 0.97 0.99 3.22 1.63 23.39 75.16
+XL-Instruct (best, 32) 18.19 52.12 31.64 8.34 5.79 1.59 5.79 1.83 38.44 76.72
Aya Expanse 8B 29.90 58.21 56.91 56.68 1.11 1.02 3.04 2.94 59.29 76.26
+XL-Instruct (best, 32) 32.31 63.24 59.53 63.22 2.21 0.78 5.85 3.66 60.01 77.70
Table 3: Win Rates of LLMs and their XL-Instruct fine-tuned counterparts on m-AlpacaEval against GPT-4o-Mini, with GPT-4o as the judge. For each model, we choose the best cross-lingual performing baseline from Figure 2 and evaluate transfer on m-AlpacaEval. Consistent improvement across all models and pairs shows strong zero-shot transfer from cross-lingual tuning, for both multilingual and English-only generation. Best scores are bolded and cells are highlighted proportionate to performance gain.
Having seen task-specific improvements, we now seek to evaluate the performance of models fine-tuned with XL-Instruct on multilingual and English open-ended generation, since these are arguably the more common use cases of LLMs. For this purpose, we first construct the m-AlpacaEval benchmark by machine translating the AlpacaEval test set into our 8 languages of interest, following related efforts to create m-ArenaHard (Dang et al., 2024). We use GPT-4o for translation of the prompts. The evaluation setup is similar to XL-AlpacaEval, wherein GPT-4o-Mini is the reference model and GPT-4o is the judge.
We show the results in Table 3, using the base and instruct models from Figure 2 along with the best-performing cross-lingual baselines. We observe significant and consistent zero-shot transfer across all models and languages. For multilingual generation, the gains are strongest for the languages a model is pretrained on, similar to our observations for cross-lingual generation. This is particularly evident in the Qwen and Aya models. EuroLLM Instruct, on the other hand, achieves stable performances across all languages and relatively strongest win rates for the lower-resourced languages. Interestingly, one can also see consistent gains in English-only generation, despite there being no English responses on the target side! This suggests that all of these models, which are sufficiently proficient in English, can learn improved response structure and formatting from the cross-lingual responses. These results are quite encouraging, since they suggest cross-lingual fine-tuning need not come at the cost of standard ‘monolingual’ generation performance – in fact, it can result in further boosts.
7 Conclusion
In this work, we propose data resources for advancing cross-lingual open-ended generation—loosely defined as a task in which the query and the desired (open-ended) response are in different languages. This can be viewed as a distinct yet crucial subtask of multilingual generation. While cross-lingual generation may also include more complex scenarios, such as providing context in one language while the query and response are in another (or even multiple) languages, we focus here on the simpler scenario: queries posed in English with responses required in one of eight target languages – which includes high, medium, and low-resource EU and non-EU languages.
With this goal in mind, we make three key contributions. First, we introduce the XL-AlpacaEval benchmark to evaluate the current state of open LLMs, and report poor performances and significant gaps against GPT-4o-Mini. Second, we propose the XL-Instruct technique, and show that this synthetic data can substantially boost cross-lingual performance, both in terms of win rates and fine-grained quality metrics. Third, we show that it exhibits strong zero-shot transfer to monolingual generation, both in English and beyond. Based on these results, we strongly encourage researchers to post-train their multilingual LLMs with XL-Instruct data, and shall publicly release our versions of XL-Instruct and XL-AlpacaEval datasets to support this.
8 Ethics Statement
There has been some concern in the literature that iterative training on synthetic data could eventually lead to model collapse (Shumailov et al., 2024). Like any other synthetic data technique, XL-Instruct could also share similar risks, especially since its seed data is sourced from the Web. We conduct fine-grained evaluation along various quality metrics, including Objectivity and Naturalness, to ensure that our fine-tuned models continue producing neutral, unbiased, and natural-sounding outputs and observe performance comparable to a fully instruction-tuned model fine-tuned on both human and synthetic data – which helps alleviate such concerns. We note that later works have also shown that mixing synthetic with human-generated data could avoid model collapse (Seddik et al., 2024; Gerstgrasser et al., 2024). While outside the scope of this work, it would be interesting to explore this angle as part of future research, and observe how XL-Instruct interacts with data from other tasks
Appendix A The XL-AlpacaEval Benchmark
Here we provide some additional details on the creation of the XL-AlpacaEval benchmark, which has 797 cross-lingual prompts in total, and currently supports 11 languages - the 8 languages used for the primary experiments in this work (Chinese, German, Hindi, Hungarian, Irish, Lithuanian, Maltese and Portuguese) and 3 additional languages (French, Finnish and Turkish) which we use for zero-shot evaluation in future sections. It is trivial to extend it to other languages – one simply has to run a script to append cross-lingual generation instructions (Section A.2) to our filtered AlpacaEval test set (Section A.1) and such extensions are being planned as a part of future work.
A.1 Manual Verification
Prompt ID Prompt Text
183 Write a story about Anakin Skywalker encountering a Jedi who speaks and acts like a 1920s British aristocrat.
200 Write ”Test”
350 I’m an English speaker trying to learn Japanese Kanji using mnemonics. Mnemonics for Kanji are created from the primitives that make them up. The Kanji for Tax has the primitives wheat and devil, so an example would be, ”Taxes are like the devil taking away your hard earned wheat”. Can you create a mnemonic for the Kanji meaning Wish that has the primitives clock and heart?
458 Give me a list of 5 words where the letters of the words are in alphabetical order. One example: ”doors”. ”d” comes before ”o”, ”o” comes before ”r”, and ”r” comes before ”s”.
476 Rewrite the given text and correct grammar, spelling, and punctuation errors. If you’d told me year ago that today I would finish a marathon, I would of laughed. Your support had a huge affect on me!
495 During writing, we added an asterisk for the word that did not come to mind. You will need to provide several examples to demonstrate all the words that can be used in the sentence instead of the asterisk.
635 Correct the transcription of an excerpt containing errors. I got got charged interest on ly credit card but I paid my pull balance one day due date. I not missed a pavement year yet. Man you reverse the interest charge?
662 You should capitalize the sentence according to the guide. Guide: Every other letter alternates between lower case and upper case. Sentence: A giant spider blocks your path.
663 Create alliterations by finding synonyms for words in the given sentence. David wears a hat everyday.
714 Rewrite the text and correct the spelling errors. It solves problems comon and uniqe to every team.
Table 4: Culturally specific prompts removed from the AlpacaEval dataset.
        Prompts
        Answer in {} language         Output an answer in {} language
        Generate your answer in {} language         Respond in {} language
        Produce an answer in {} language         Please write in {} language
Table 5: Cross-Lingual Generation Instructions
Before creating our cross-lingual benchmark, we conduct a rigorous stage of manual verification to ensure that the prompts are suitable for answering cross-lingually. In Table 4, we show the prompts we removed from AlpacaEval that were too culturally specific (for instance, prompt 183) or tailored towards eliciting an English response (prompts 350 and 714). In the latter, we felt mandating a non-English response might make evaluating a “correct” response challenging. In other cases where the prompt simply requested a response in English, we replaced with a generic templated variable {language} for downstream substitution with the name of the desired target language. This leaves us with a total of 797 prompts. It is important to note that as far as possible, we tried to keep complex multi-step, multilingual prompts in our evaluation set, and only removed cases that were clearly invalid – in keeping with the goal of this work to build robust cross-lingual models.
A.2 Generation prompts
Next, we randomly sample prompts from a list of cross-lingual generation instructions (given in Table 5), and append it to each prompt in the filtered test set from the previous stage. To add further diversity to the instructions in the benchmark, we remove the word “language” from the prompts given in Table 5 – thus converting “Answer in German language” to “Answer in German”. This leads to the creation of the final XL-AlpacaEval benchmark.
Appendix B Experiments
B.1 XL-AlpacaEval Results
Model Avg zho deu hin hun gle lit mlt por
EuroLLM 9B 7.36 8.97 9.96 4.49 4.13 6.09 9.94 4.66 10.61
+2K instructions 18.63 18.77 23.65 13.22 13.70 16.03 25.48 14.75 23.47
+8K instructions 21.54 20.98 26.76 16.26 17.27 20.99 28.52 15.72 25.81
+32K instructions 20.54 21.24 24.07 15.26 17.11 21.08 28.09 15.64 21.79
EuroLLM 9B Instruct 12.70 14.82 16.49 8.23 8.66 9.37 16.57 8.51 18.94
+32 instructions 20.84 23.52 22.96 13.10 17.37 17.10 25.61 21.30 25.79
+256 instructions 17.83 21.13 21.73 12.90 14.05 13.25 21.13 15.32 23.11
+2K instructions 21.18 23.62 24.39 14.49 16.63 20.17 27.87 18.02 24.22
+8K instructions 19.75 23.10 22.65 14.50 14.97 20.55 26.92 15.34 19.96
Qwen 2.5 7B 5.80 12.62 6.36 3.40 2.73 4.50 4.33 2.62 9.82
+2K instructions 13.85 33.64 18.37 6.67 6.50 5.00 10.73 3.63 26.22
+8K instructions 13.91 34.22 19.80 6.61 6.28 3.92 10.22 3.07 27.13
+32K instructions 13.94 34.29 18.88 6.88 5.72 5.44 10.77 3.36 26.18
Qwen 2.5 7B Instruct 16.73 44.63 16.35 9.59 6.82 7.17 14.68 3.69 30.88
+32 instructions 22.85 50.16 31.66 12.36 12.52 8.66 19.40 4.91 43.10
+256 instructions 17.00 38.04 22.45 9.46 7.85 5.39 15.86 4.02 32.92
+2K instructions 14.97 36.17 18.95 8.44 7.14 5.06 12.02 3.02 28.92
+8K instructions 15.57 42.74 18.85 8.32 6.54 4.41 11.99 3.49 28.19
Aya Expanse 8B 35.67 57.22 60.27 56.99 8.62 10.43 19.54 9.51 62.75
+32 instructions 38.61 64.08 65.07 59.76 10.71 11.72 21.57 10.70 65.28
+256 instructions 30.39 55.65 52.93 44.50 6.51 6.45 17.90 6.07 53.10
+2K instructions 25.30 41.84 46.43 37.03 6.77 4.55 15.94 3.75 46.12
+8K instructions 23.32 43.16 42.23 28.19 5.44 6.00 15.94 4.91 40.72
Table 6: Full language-wise Win Rates against GPT-4o-Mini on XL-AlpacaEval, after LoRA fine-tuning on varying sizes of XL-Instruct data on different LLMs. GPT-4o is the judge. The best scores per model are highlighted in bold.
Full Results
In Table 6, we show the complete language-wise results for each base and instruct model we tuned on varying sizes of XL-Instruct data. Models like EuroLLM and Qwen continue improving until 8K-32K instructions, with gains diminishing in the last 24K instructions. This is likely because we sort the instructions in order of translation quality, and sample them accordingly, reducing the gains. It is possible that improving the translation quality further could result in larger gains. For preference-optimized (PO’ed) instruction-tuned models, performance saturates at 32 instructions, and 2K instructions with non-PO’ed models like EuroLLM 9B Instruct. The largest gains across all models are consistently for the languages included during pretraining – for instance, Qwen 7B improves on Chinese win rates from 12.62 to 34.29 and in Portuguese from 9.82 to 27.13, suggesting the criticality of this stage in building multilingual LLMs.
Model Avg French Finnish Turkish
EuroLLM 9B 7.80 9.69 9.78 3.94
EuroLLM 9B Instruct 14.08 19.39 14.05 8.81
EuroLLM 9B XL-Instruct (best) 20.62 25.8 22.72 13.33
Table 7: Fine-tuning with XL-Instruct yields zero-shot boosts in cross-lingual performance. Scores represent zero-shot win rates of various LLMs against GPT-4o-Mini, with GPT-4o as a judge. For the XL-Instruct baseline, we use the best-performing model from Figure 2.
Zero-Shot Results
We show in Table 7 evaluations on zero-shot performance after fine-tuning with XL-Instruct. We choose French, Finnish and Turkish, 3 languages the EuroLLM model is pre-trained on, and observe huge gains in win rates, largely outperforming even the EuroLLM 9B Instruct model. This shows that even if done only for a few languages, XL-Instruct can still result in significant transfer that improves performance in others. We hypothesize that this is likely because the model is able to learn formatting, response structure etc. from this process, which also supports the boosts in English generation one observes in Table 3.
Fine-grained Evaluation
Finally, we also tabulate the macro-averaged GPT-4o scorings on Precision, Informativeness, Naturalness and Objectivity metrics (evaluation prompts and rubrics available at https://tinyurl.com/xl-gen-eval). As noted previously, the EuroLLM 9B model performs the worst, but the XL-Instruct model performs comparably or slightly better than EuroLLM 9B Instruct, indicating the efficacy of our data generation method.
Model Average Precision Informativeness Naturalness Objectivity
EuroLLM 9B 2.43 2.52 2.69 2.25 2.27
EuroLLM 9B Instruct 3.56 3.68 3.80 3.54 3.23
EuroLLM 9B XL-Instruct (best) 3.60 3.63 3.88 3.64 3.24
Table 8: Performance of EuroLLM 9B models evaluated on Precision, Informativeness, Naturalness, and Objectivity, with the average of these metrics.
B.2 Comparison with X-Instruction
Model Avg Finnish Hindi Turkish
EuroLLM 9B + X-Instruction (full 1M) 9.73 14.35 7.22 7.63
EuroLLM 9B + X-Instruction (40K) 10.44 13.69 8.76 8.86
EuroLLM 9B + XL-Instruct (naive, 40K) 12.06 15.3 10.9 9.98
EuroLLM 9B + XL-Instruct (best of 3, 40K) 17.82 23.15 15.8 14.52
Table 9: Performances of EuroLLM 9B models fine-tuned on X-Instruction and XL-Instruct data
We also compare the efficacy of our method with the most similar work to ours and the only cross-lingual open-ended generation dataset we are aware of: X-Instruction (Li et al., 2024a). In this work, the authors prompt a teacher LLM for reverse instructions directly in the low-resource language. Given this initially results in instructions of much poorer quality due to poor teacher capabilites, they also conduct iterative scoring and refinement to improve quality, achieving impressive results. They also release their dataset publicly at https://huggingface.co/datasets/James-WYang/X-Instruction.
We use the Hindi, Finnish, and Turkish splits of this dataset since these are supported by EuroLLM and are also available in X-Instruction. We also generate XL-Instruct data in these languages, by redoing the XL-Instruct pipeline (Section 4) from Stage 3 (Response Translation) for these languages. We LoRA fine-tune EuroLLM 9B on various X-Instruction and XL-Instruct datasets. For the former, we use both the entire 1M sized dataset available for these languages (in total), and a 40K instructions subset which is more comparable to our XL-Instruct baselines. For XL-Instruct, we train two baselines – one trained on ‘naive’ translations (ie. using only EuroLLM 9B Instruct) and another using the ‘best-of-3’ method (please refer Section 4.3 for a detailed explanation).
We see that both XL-Instruct baselines significantly outperform X-Instruction, with our best model achieving a 70.68% improvement over the latter – showcasing the relative superiority of our pipeline. This also suggests it might be more effective to prompt a teacher model in English due to inherently superior capabilities, and we hypothesize it might allow for greater quality and diversity in responses, as well as allowing for more complex operations – like refinement following specifically defined, custom criteria.
B.3 Ablations
Lastly, we conduct an ablation to verify the importance of the translation selection strategy. Given the cross-lingual part of the dataset mainly comes from Machine Translations, and translations can be quite noisy, we experiment with 2 MT techniques, ‘naive’ and ‘best-of-3’ responses. We also include a ‘random’ sampling strategy, where random responses are chosen for subsampling, regardless of MT quality. We fine-tune the EuroLLM 9B and EuroLLM 9B Instruct models using 8K and 32 instructions respectively, which are respectively the optimal SFT data sizes for each model (check Figure 2).
For the instruct model, ‘best of 3’ introduces significant improvements over naive or random sampling strategies, taking the average win rate from 18.55 to 20.84. This is likely because at the tiny scale of 32 instructions, target response quality matters hugely and significantly impacts performance. For EuroLLM 9B, which is fine-tuned on 8K instructions, performance still improves for most languages with the best-of-3 technique. The only cases where it drops are for the least-resourced languages like Irish and Maltese, which makes the average score much lower. It is possible the CometKiwi model we use for Quality Estimation is not very well-suited for such low-resource languages. As a result, we hypothesize that best-of-3 might sometimes end up choosing a worse translation than the naive method – which uses EuroLLM, a model known to have strong MT capabilities for all these languages.
Model Avg zho deu hin hun gle lit mlt por
EuroLLM 9B 7.36 8.97 9.96 4.49 4.13 6.09 9.94 4.66 10.61
+8K instructions (random) 22.69 22.66 25.71 15.59 18.45 22.40 29.52 20.50 26.72
+8K instructions (naive) 21.17 20.26 24.63 15.53 16.68 21.96 28.33 18.24 23.75
+8K instructions (best of 3) 21.54 20.98 26.76 16.26 17.27 20.99 28.52 15.72 25.81
EuroLLM 9B Instruct 12.70 14.82 16.49 8.23 8.66 9.37 16.57 8.51 18.94
+32 instructions (random) 18.49 22.21 22.69 12.70 15.18 15.35 21.87 14.12 23.79
+32 instructions (naive) 18.55 22.20 20.16 12.18 13.70 15.14 23.64 17.55 23.84
+32 instructions (best of 3) 20.84 23.52 22.96 13.10 17.37 17.10 25.61 21.30 25.79
Table 10: Ablations of the strategy for selecting response translations for the EuroLLM 9B and EuroLLM 9B Instruct models.